Q 1.16: 6.0 x 106 m + 2.0 x 105 m= A 6.2 x 105 m B 8.0 x 106 m C 8.0 x 105 m D 6.2 x 106 m
Answer #1

here,

a = 6 * 10^6 m

b = 2 * 10^5 m

b = 0.2 * 10^6 m

the resultant , R = a + b

R = 6 * 10^6 m + 0.2 * 10^6 m

R = 6.2 * 10^6 m

the correct option is (D) 6.2 * 10^6 m